OpenStudy (anonymous):

Car speed. On a sales trip, Mark drives the 600 mi to Richmond averaging a certain speed. The return trip is made at an average speed that is 10 mph slower. Total time for the round trip is 22 hr. Find Mark’s average speed on each part of the trip.

OpenStudy (saifoo.khan):

\[Speed = \frac{distance}{time}\]

OpenStudy (anonymous):

27 ?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (anonymous):

54.5 miles per hour

OpenStudy (saifoo.khan):

Make the data first, To richmond, time = 600/x from Richmond, 600 /(x-10) now equate, \[\frac{600}{x} - \frac{600}{x-10} = 22\]Can you solve now?

OpenStudy (anonymous):

you lost me.

OpenStudy (saifoo.khan):

Where?

OpenStudy (anonymous):

oh i see I complety left out the 10 mph, sorry

OpenStudy (saifoo.khan):

Np, so you get it now?

OpenStudy (anonymous):

yes lol,

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

your awesome

OpenStudy (saifoo.khan):

Please hit good answer. And thanks.

OpenStudy (saifoo.khan):

Tx